OpenStudy (anonymous):

by removing the absolute values, express (4-x^2)/|x+2| as a peicewise-defined function

OpenStudy (anonymous):

\[f(x) = |x+2| = \left\{\begin{array}{rcc} -x-2 & \text{if} & x <2\\ x+2& \text{if} & x>2 \end{array} \right. \]is a start

OpenStudy (anonymous):

damn all that formatting and it is wrong!

OpenStudy (anonymous):

the restrictions wouldn't be x< -2 and x> -2

OpenStudy (anonymous):

\[f(x) = |x+2| = \left\{\begin{array}{rcc} -x-2 & \text{if} & x <-2\\ x+2& \text{if} & x>-2 \end{array} \right.\]

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

yes but the answer choices are x + 2, x > 2, −(x + 2), x <2 2. f(x) = x−2, x > −2, 2−x, x < −2. 3. f(x) = 2−x, x > −2, x−2, x < −2. 4. f(x) = x + 2, x > −2, −(x + 2), x < −2. 5. f(x) = x−2, x > 2, 2−x, x < 2. 6. f(x) = −(x + 2), x > 2, x + 2, x < 2.

OpenStudy (anonymous):

fine i just did the denominator

OpenStudy (anonymous):

factor the numerator if \(x>-2\) there will be a common factor of \(x+2\) top and bottom cancel it

OpenStudy (anonymous):

if \(x<-2\) there will be a factor of \(x+2\) in the top and \(-x-2\) in the bottom and \[\frac{x+2}{-x-2}=-1\]

OpenStudy (anonymous):

I did the top which would simplify It -(x-2)(x+2)/(x+2) so the then the x+2 cancels out

OpenStudy (anonymous):

im confused as to what your saying

OpenStudy (anonymous):

top is \((2-x)(2+x)\) right?

OpenStudy (anonymous):

if \(x>-2\) you can cancel the \(2+x\) top and bottom and get just \(2-x\)

OpenStudy (anonymous):

whereas if \(x<-2\) the denominator is \(-x-2\)

OpenStudy (anonymous):

you do not have a factor of \(-x-2\) in the numerator, however you can still cancel because \[\frac{2+x}{-x-2}=-1\] leaving you with \(-(2-x)\) or \(x-2\)

OpenStudy (anonymous):

I get it so (2-x) x>-2 -x-2 x<-2